Q4. A nationwide supermarket chain has launched a new loyalty card system. Customers obtain 1 point for each £10 pounds they spend instore or online throughout the year. Points are redeemed as vouchers issued at a rate £1 per 10 points. The number of pounds that a customer spends at this supermarket chain across a year is Normally distributed with a mean of £5000 and standard deviation of £1250. (a) What is the probability that a customer spends more than £5000 in any given year? (b) What percentage of customers receive more than £75 in vouchers each year? (c) What percentage of customers receive between £40 and £75 in vouchers each year? (d) The marketing department want to advertise this loyalty card by saying that "90% of loyalty card holders receive at least £X in vouchers every year". What is the highest number that could be truthfully) used in this statement?
Answer #1

Answer:-

Given that:-

Let,X: Number of pounds spent by the customer at the supermarket chain.

X • N = £5000,0”. £12502)

also, 1point for \pounds 10. spent.

and £1 are redeemed as voucher for 10 points

a) The probability that a customer spends more than £5000 in any given year can be calculated as

P(X > £5000) = P(X-M £5000 - £12500 o by central limit theorem.

  P(Z>\frac{\pounds 5000-\pounds 5000}{\pounds 12500})

=P(Z>0)

P(X>\pounds 5000)=0.5

b) A customer receives suppose \pounds k in voucher that means customer earned 10*\pounds k points that also means he/she spent 10*10*\pounds k

\thereforeThe percentage of customers receive more than \pounds 75 in vouchers each year can be calculated as P(X>\pounds 7500)*100

Hence, P(X>\pounds 7500)

  =P(\frac{X-\mu}{\sigma }>\frac{\pounds 7500-\mu}{\sigma }) by CLT

=P(Z>\frac{\pounds 7500-\pounds 7500}{\pounds 1250 })

=P(Z>2)

P(X>\pounds 7500)=0.02275

\therefore P(X>\pounds 7500)*100=2.275\%

\therefore 2.275\% of the customers receive more than \pounds 75 in vouchers each year.

c) The percentage of customers receive between \pounds 40 and \pounds 75 in vouchers each year can be calculated as the same way .

P(\pounds 4000<X<\pounds 7500)*100

\therefore P(\pounds 4000<X<\pounds 7500)

  =P(\frac{\pounds 4000-\mu}{\sigma }<Z<\frac{\pounds 7500-\mu}{\sigma }) by CLT

=P(\frac{\pounds 4000-\pounds 5000}{\pounds 1250 }<Z<\frac{\pounds 7500-\pounds 5000}{\pounds 1250 })=P(-0.8<Z<2)

\therefore P(\pounds 4000<X<\pounds 7500)=0.76539\therefore P(\pounds 4000<X<\pounds 7500)*100=76.539\%\therefore76.539\% of the customers receive between \pounds 40 and \pounds 75 in each year.

d) 90% of loyalty card holders receive at least \pounds X in vouchers each year. \pounds X can be caluclated as below.

Let , for the sake or convenience we replace our random variable x as Y.

P(Y>\pounds X)=0.90

\therefore P(\frac{y-\mu}{\sigma }> \frac{\pounds X-\mu}{\sigma })=0.90 by CLT.

\therefore P(Z>\frac{\pounds X-\pounds 5000}{\pounds 1250})=0.90

\thereforeBy using statistic table we get z as z= -1.282

\therefore \frac{\pounds X-\pounds 5000}{\pounds 1250}=-1.282

\therefore \pounds X=\pounds 3397.5

90% loyalty card holders will receive at least \pounds 33.975 every year in voucher.